OpenStudy (anonymous):

perform the indicated operations & simplify the result. Leave the answer in factored form. ((x/x+4) + 1) / ((12/x^2-16) + 1)

OpenStudy (anonymous):

It might be easier to see like this: \[\frac{\frac{x}{x+4}+1}{\frac{12}{x^{2}-16}+1}\]

OpenStudy (anonymous):

yeah, I wasn't sure how to type it in like that :/

OpenStudy (anonymous):

No worries.

hero (hero):

Actually, I'd probably write it a differnt way initially.

OpenStudy (anonymous):

i wouldn't add

hero (hero):

\[\left(\frac{x}{x+4}+1\right) \div\ \left(\frac{12}{x^2-16} + 1\right)\]

hero (hero):

There. That's much better

hero (hero):

Now add, then divide

OpenStudy (anonymous):

i would start with \[\frac{\frac{x}{x+4}+1}{\frac{12}{(x-4)(x+4)}+1}\] and multiply numerator and denominator by \((x+4)(x-4)\) cancelling merrily as i went along

hero (hero):

I'm trying to keep it simple @satellite73

OpenStudy (anonymous):

\[\frac{\frac{x}{x+4}+1}{\frac{12}{(x-4)(x+4)}+1}\times \frac{(x+4)(x-4)}{(x+4)(x-4)}\] i think this is simpler, but that is of course a matter of taste

OpenStudy (anonymous):

that makes so much sense @satellite73 !! I didn't even think to factor the other denominator! been working at this package for far too long

OpenStudy (anonymous):

get rid of the compound fraction in one step, then multiply out, combine like terms etc

OpenStudy (anonymous):

you get \[\frac{x(x+4)+(x+4)(x-4)}{12+(x+4)(x-4)}\] and then you can do the algebra top and bottom

hero (hero):

\[\left(\frac{x}{x+4}+\frac{x+4}{x+4}\right) \div\ \left(\frac{12}{x^2 - 16} + \frac{x^2 - 16}{x^2 - 16}\right)\]

OpenStudy (anonymous):

I was trying to multiply everything by \[x^{2} - 16\] ..silly girl

hero (hero):

\[\left(\frac{2x + 4}{x+4}\right) \div\ \left(\frac{x^2 -4}{x^2 - 16} \right)\]

hero (hero):

\[\left(\frac{2x + 4}{x+4}\right) \times\ \left(\frac{x^2 -16}{x^2 - 4} \right)\]

hero (hero):

And you can still cancel from there

hero (hero):

My method isn't too bad either

hero (hero):

I wish I could give myself a medal

OpenStudy (anonymous):

@refusetofail not silly at all that is what i did, it is just easier to see if when it is in factored form

OpenStudy (anonymous):

so.. wait, I'm kind of confused how to get the actual answer through the simplification..